# A Practitioner’s Guide to Kolmogorov–Arnold Networks

### v2

@@ Abstract
The so-called Kolmogorov–Arnold Networks (KANs), whose design is merely inspired, rather than dictated, by the Kolmogorov superposition theorem, have emerged as a promising alternative to traditional Multilayer Perceptrons (MLPs). This review provides a systematic and comprehensive overview of the rapidly expanding KAN landscape. By collecting and categorizing a large set of open-source implementations, we map the vibrant ecosystem supporting modern KAN development. We organize the review around four core themes:
(i) presenting a precise history of Kolmogorov’s superposition theory toward neural-network formulations (Sec. 3);
(ii) establishing the formal equivalence between KANs and MLPs (Secs. 4,5);
(iii) analyzing the critical role of basis functions (Sec. 6); and
(iv) organizing recent advancements in accuracy, efficiency, regularization, and convergence (Secs. 7, 8, 9, 10).
Finally, we provide a practical Choose-Your-KAN guide (Sec. 11) to assist practitioners in selecting appropriate architectures, and we close by identifying current research gaps and future directions (Sec. 12). The associated GitHub repository (https://github.com/AmirNoori68/kan-review) complements this paper and serves as a structured reference for ongoing KAN research.
@@ 1 Introduction
Multilayer perceptrons are universal approximators and remain a standard building block for regression, function approximation, and pattern recognition across scientific and engineering applications. A major extension of this paradigm is the physics-informed neural network (PINN) framework of Raissi et al. [1], which has rapidly grown and enabled large-scale studies.
Since then, numerous advances have broadened the scope of MLP-based PINNs. Fractional operators were addressed by fPINNs [2], while uncertainty quantification for forward and inverse problems was introduced in [3]. Extensions such as VPINNs [4], XPINNs for domain decomposition [5], and multi-fidelity PINNs [6] improved accuracy and scalability. The framework also inspired convergence analyses [7], NTK-based diagnostics [8], and practical toolkits like DeepXDE [9]. More recent contributions include multi-stage training with near machine-precision accuracy [10], adaptive residual weighting [11], and distributed solvers for extreme-scale PDEs [12, 13, 14, 15]. Collectively, these developments establish MLP–PINNs as a mature reference standard in scientific machine learning, many of which, as we show in Section 5, have directly inspired KAN development.
@@
KANs [26] replace fixed activations with learnable univariate functions on edges, motivated by the Kolmogorov representation theorem. The original model uses B-spline bases [26], but the framework is modular: Chebyshev polynomials, Gaussian kernels, and other basis families can be substituted as needed. This flexibility is central, since the chosen basis controls smoothness, locality, and spectral behavior, directly affecting expressivity and interpretability. The combination of adaptive basis learning and interchangeable basis families is therefore a core advantage of KANs. KANs also include a growing set of architectural variants, optimization strategies, and theoretical developments.
A number of recent surveys have appeared on KANs, each highlighting different facets of this emerging architecture. Notable contributions include those by Andrade et al. [27], Basina et al. [28], Beatrize et al. [29], Faroughi et al. [30], Kilani et al. [31], Hou et al. [32], Dutta et al. [33], Essahraui et al. [34], and Somvanshi et al. [35], which together provide valuable entry points into the literature. Our review builds on these efforts by seeking to offer a more systematic and comprehensive perspective: rather than cataloguing studies in isolation, we integrate theoretical, architectural, optimization, and application viewpoints into a structured roadmap. By combining comparative analysis, methodological insights, and practical guidance, our goal is to provide a resource that complements existing surveys and serves readers aiming to both understand and apply KANs. Table 1 presents a structured roadmap of the paper’s sections and themes.
Table 1: Paper roadmap (clickable links to sections).
1 Introduction 2 MLP Basics 3 KST Toward Neural Networks 4 Bridging KANs and MLPs 5 How KANs Extend the MLPs Landscape 6 Basis Functions $\triangleright$ 6.1 B-spline $\triangleright$ 6.2 Chebyshev Polynomial $\triangleright$ 6.3 ReLU $\triangleright$ 6.4 Jacobi and General Polynomials $\triangleright$ 6.5 Gaussian RBF $\triangleright$ 6.6 Fourier $\triangleright$ 6.7 Wavelet $\triangleright$ 6.8 Finite-Basis $\triangleright$ 6.9 SincKAN 7 Accuracy Improvement $\triangleright$ 7.1 Physics Constraints & Loss Design $\triangleright$ 7.2 Adaptive Sampling & Grids $\triangleright$ 7.3 Domain Decomposition $\triangleright$ 7.4 Function Decomposition $\triangleright$ 7.5 Hybrid/Ensemble & Data $\triangleright$ 7.6 Sequence/Attention Hybrids $\triangleright$ 7.7 Discontinuities & Sharp Gradients $\triangleright$ 7.8 Optimization & Adaptive Training 8 Efficiency Improvement $\triangleright$ 8.1 Parallelism, GPU, and JAX Engineering $\triangleright$ 8.2 Matrix Optimization & Efficient Bases 9 Sparsity & Regularization 10 Scaling Laws & Convergence $\triangleright$ 10.1 Theoretical Approximation Rates $\triangleright$ 10.2 Spectral Bias $\triangleright$ 10.3 NTK-Based Convergence $\triangleright$ 10.4 Practical Trade-offs 11 Practical “Choose–Your–KAN” Guide 12 Current Gaps and Path Forward 13 Conclusion
@@
When it comes to deciding whether MLP-PINNs or KAN-PIKANs are the better choice, the comparison is far from straightforward, with studies that emphasize fairness often reaching divergent or contradictory conclusions [36] [^1] https://github.com/yu-rp/KANbeFair, [37, 38]. A key reason is that KANs are not a monolithic architecture; their performance is critically influenced by the choice of basis function. For example, a comprehensive comparison by Farea and Celebi [39] [^2] https://github.com/afrah/pinn_learnable_activation found that the optimal basis (e.g., B-spline, Fourier, Gaussian) varies significantly with the PDE being solved, making a simple “KAN vs. MLP” verdict insufficient.
@@
Despite these complexities, Table 2 summarizes reported outcomes across regression, function approximation, and PDE-solving tasks. The entries reflect prevailing tendencies in the literature rather than definitive results for every case. A consistent pattern is that KANs—particularly when equipped with specialized basis functions—tend to match or outperform vanilla MLPs and PINNs in terms of accuracy and convergence speed.
@@
Table 2: Representative performance trends comparing KAN-based models with MLPs. Results are aggregated from the literature and reflect general outcomes. “Slower training” refers to higher per-iteration cost, though total wall-clock time can be different due to fewer epochs to convergence. All rows compare KANs to plain MLP/PINN baselines without advanced features.
Ref. Accuracy Convergence/Time(per-iter.) Basis Functions
@@
KANs are rapidly expanding on open-access platforms, with diverse implementations and variants shared by the community. Curated repositories such as [75] [^3] https://github.com/mintisan/awesome-kan track the latest developments across different fields, making this growth more accessible to readers. Table LABEL:githubs summarizes representative GitHub repositories that illustrate this fast-growing ecosystem in regression, function approximation, and PDE solving. These resources are also referenced in footnotes throughout this study when further details of specific methods are discussed.
Organization (Table 1). Section 2 reviews MLP basics. Section 3 introduces Kolmogorov Superposition Theory and its refinements. Section 4 connects KANs to classical neural networks. Section 6 surveys basis families and computational trade-offs. Section 7 covers accuracy-improvement strategies, and Section 8 addresses efficiency via parallelism/GPU/JAX and matrix optimization. Section 9 treats sparsity, regularization, and Bayesian variants, while Section 10 discusses scaling laws and convergence. Section 11 presents the practical “Choose–Your–KAN” guide. We conclude with final remarks in Section 13.
2 MLP Basics
$\mathbf{z}^{(0)}$ $=\mathbf{x},$ (1)
where
  • $n_{\ell}$ is the width (number of neurons) of layer $\ell$,
3 KST Toward Neural Networks
Superpositions and Hilbert’s Question. A superposition is any function obtained by repeatedly composing simpler functions. Lorentz [82] illustrates this idea with the example
@@ 3 KST Toward Neural Networks
for scalar inputs $x,y,z\in\mathbb{R}$, showing how a multivariate mapping can be built from bivariate ones. Even the elementary identity
@@
is a superposition of binary additions.
Hilbert’s 13th Problem [76], posed at the 1900 International Congress of Mathematicians, asked whether every continuous function of three variables
$f:[0,1]^{3}\to\mathbb{R}$
can be represented as a finite superposition of continuous bivariate functions. Hilbert conjectured that such a representation was impossible. However, the later work of Kolmogorov and Arnol’d overturned this view, showing that every continuous function of $s$ variables can be expressed using only sums and superpositions of univariate functions—a result that fundamentally reshaped the theory of functional representation.
Kolmogorov and Arnol’d: resolving the 13th Problem. Kolmogorov’s 1956 note [77] introduced the idea that multivariate functions might be representable via functions of fewer variables. Arnol’d, his student, confirmed this for the three–variable case in 1957 [78], giving an explicit decomposition of a general $f(x,y,z)$ into superpositions of continuous bivariate functions. Later that year, Kolmogorov published his full and elementary solution [79]: any continuous function of $n>2$ variables can be written using only univariate functions and addition. Thus Hilbert’s conjecture was disproved.
The Kolmogorov Superposition Theorem (KST). Let $E_{1}=[0,1]$ and $E_{n}=[0,1]^{n}$. Kolmogorov proved that for every $n>2$ there exist continuous universal inner functions
@@
independent of the target function $f$, such that every $f\in C(E_{n})$ admits the representation
@@
where each $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ is continuous and depends on $f$, while the inner functions $\phi_{q,p}$ are universal. KST shows that multivariate continuity can be synthesized entirely from univariate nonlinearities, a structural insight that later motivated two–layer interpretations and neural–network analogies.
@@
Refinements to KST. Kolmogorov’s constructive formulation (4) established the universality of the inner functions $\phi_{q,p}$, but it did so at considerable structural cost: the representation requires $2n^{2}{+}n$ distinct inner functions and 2n+12n{+}1 outer functions. While theoretically elegant, this level of redundancy makes the original form far too complex for computation or numerical approximation. The decades following Kolmogorov’s work therefore focused on two goals: (i) reducing the number of unique functions needed, and (ii) improving their regularity. These efforts produced the modern, compressed form of KST that underlies computational constructions and later neural interpretations.
(i) Reducing the number of functions. Sternfeld [85] proved that the number of summands $2n+1$ in the Kolmogorov superposition formula is optimal and cannot be reduced. However, his functional–analytic framework shows that these 2n+12n{+}1 summands do not require 2n+12n{+}1 different inner or outer functions: the same small collection of functions can be reused repeatedly. Thus the number of distinct functions can be much smaller, even though the number of summands is fixed.
Lorentz’ reduction of outer functions. Lorentz’s 1962 exposition [82] clarified Kolmogorov’s argument and showed that the 2n+12n{+}1 outer functions $\Phi_{q}$ can be replaced by a single continuous function $\Phi$. In the bivariate case ($s=2$), with inputs $(x_{1},x_{2})\in[0,1]^{2}$, he derived the canonical representation
$f(x_{1},x_{2})=\sum_{q=1}^{5}\Phi\!\bigl(\phi_{1,q}(x_{1})+\phi_{2,q}(x_{2})\bigr),$
where the ten inner functions $\phi_{1,q}$ and $\phi_{2,q}$ are universal, but the outer function $\Phi$ depends on $f$. More generally, for $s$ variables he obtained
@@
Sprecher’s reduction of inner functions. Sprecher [83] achieved the corresponding simplification for the inner layer. He showed that the entire family $\{\phi_{q,p}\}$ can be generated from a single monotone universal function $\psi:E_{1}\to\mathbb{R}$, together with fixed positive weights $\{\lambda_{p}\}_{p=1}^{n}$ and uniform shifts $x\mapsto x+\varepsilon q$, where $\varepsilon>0$ is a small constant. In modern notation,
$\phi_{q,p}(x)=\lambda_{p}\,\psi(x+\varepsilon q),$ (6)
where the weights $\lambda_{p}>0$ depend only on the dimension and satisfy a rational independence condition. Sprecher also proved that $\psi$ may be chosen with an explicit Hölder–Lipschitz exponent.
@@
(ii) Improving regularity. The inner functions in Kolmogorov’s original proof are highly irregular—Hölder but non-differentiable. A function $u:E_{1}\to\mathbb{R}$ is Hölder continuous with exponent $\alpha\in(0,1]$ if
@@
Vitushkin and Henkin [86, 87, 88] proved that no universal inner family can possess $C^{1}$ (or $C^{s}$ with $s>1$) smoothness. Within the theoretical development of KST itself, Fridman [92] constructed universal inner maps in $\mathrm{Lip}^{1}$, the maximal regularity consistent with these impossibility results. Sprecher [93] further demonstrated that his compressed formulation (6) can also be realized with Lipschitz inner functions. More recently, Actor and Knepley introduced the first explicit algorithm aimed at computing Lipschitz universal inner functions, refining the earlier Hölder-type constructions of Köppen and Braun–Griebel [94, 95].
Computational KST. Algorithmic approaches fall into three categories:
@@
  • Exact constructive KST: compute the universal inner map and the outer functions exactly.
@@
  • Approximate KST (outer-only): fix the universal inner map and approximate or learn only the outer function.
@@
  • Approximate KST (inner–and–outer): approximate both inner and outer maps within a superposition-type architecture.
Exact constructions. Sprecher’s digit-based inner function and outer correction schemes were made fully computable and later corrected by Köppen and rigorously validated by Braun–Griebel, who established continuity and monotonicity of the corrected recursion [96, 97, 98, 99]. Constructive and guaranteed versions include Nakamura et al., who provide algorithmic constructions of the universal inner map $\psi$ and outer map $\chi$ with certified error bounds [101].

### v3

@@ Abstract
Kolmogorov–Arnold Networks (KANs), whose design is inspired—rather than dictated— by the Kolmogorov superposition theorem, have emerged as a structured alternative to MLPs. This review provides a systematic and comprehensive overview of the rapidly expanding KAN literature. The review is organized around three core themes: (i) clarifying the relationships between KANs and Kolmogorov superposition theory (KST), MLPs, and classical kernel methods; (ii) analyzing basis functions as a central design axis; and (iii) summarizing recent advances in accuracy, efficiency, regularization, and convergence.
Finally, we provide a practical “Choose–Your–KAN” guide and outline open research challenges and future directions. The accompanying GitHub repository (https://github.com/AmirNoori68/kan-review) serves as a structured reference for ongoing KAN research.
The following compact, clickable overview summarizes the main topics and organizational structure covered in this review.
Keywords: Kolmogorov–Arnold Networks; Kolmogorov superposition theorem; basis functions; kernel methods; neural network architectures; physics-informed learning; function approximation.
@@ 1 Introduction
MLPs. Multilayer perceptrons are universal approximators and remain a standard building block for regression, function approximation, and pattern recognition across scientific and engineering applications. A major extension of this paradigm is the physics-informed neural network (PINN) framework of Raissi et al. [1], which has rapidly grown and enabled large-scale studies.
Since then, numerous advances have broadened the scope of MLP-based PINNs. Fractional operators were addressed by fPINNs [2], while uncertainty quantification for forward and inverse problems was introduced in [3]. Extensions such as VPINNs [4], XPINNs for domain decomposition [5], and multi-fidelity PINNs [6] improved accuracy and scalability. The framework also inspired convergence analyses [7], NTK-based diagnostics [8], and practical toolkits like DeepXDE [9]. More recent contributions include multi-stage training with near machine-precision accuracy [10], adaptive residual weighting [11], and distributed solvers for extreme-scale PDEs [12, 13, 14, 15]. Collectively, these developments establish MLP–PINNs as a mature reference standard in scientific machine learning, many of which, as we show in Section 4.1, have directly inspired KAN development.
@@
KANs. Kolmogorov–Arnold Networks (KANs) in their modern multilayer form were introduced by Liu et al. [26] as a structured alternative to standard MLPs, in which fixed pointwise activations are replaced by learnable univariate functions defined on network edges and composed across layers. This design is motivated by the Kolmogorov representation theorem, but departs from its classical construction by learning the inner functions directly from data.
The original KAN architecture employs B-spline bases; however, the framework is inherently modular, allowing alternative basis families—such as Chebyshev polynomials, Gaussian kernels, Fourier features, and others—to be substituted without changing the overall architecture. This modularity is fundamental, as the choice of basis directly governs smoothness, locality, spectral content, and numerical stability, thereby influencing expressivity, interpretability, and optimization behavior. As a result, adaptive basis learning combined with interchangeable basis families constitutes a defining feature of KANs, alongside a growing collection of architectural variants, optimization strategies, and theoretical analyses.
In parallel with these methodological developments, KAN implementations have rapidly expanded across open-access platforms, with diverse variants shared by the research community. Curated resources such as [77] [^1] https://github.com/mintisan/awesome-kan provide a centralized overview of recent developments across regression, function approximation, and PDE-solving tasks. Table LABEL:githubs summarizes representative GitHub repositories illustrating this fast-growing ecosystem, which are also referenced throughout this review when discussing specific methods and design choices.
@@
MLPs vs. KANs. When it comes to deciding whether MLP-PINNs or KAN-PIKANs are the better choice, the comparison is far from straightforward, with studies that emphasize fairness often reaching divergent or contradictory conclusions [38] [^2] https://github.com/yu-rp/KANbeFair, [39, 40]. A key reason is that KANs are not a monolithic architecture; their performance is critically influenced by the choice of basis function. For example, a comprehensive comparison by Farea and Celebi [41] [^3] https://github.com/afrah/pinn_learnable_activation found that the optimal basis (e.g., B-spline, Fourier, Gaussian) varies significantly with the PDE being solved, making a simple “KAN vs. MLP” verdict insufficient.
@@
Despite these complexities, Table LABEL:KAN_vs_MLP summarizes reported outcomes across regression, function approximation, and PDE-solving tasks. The entries reflect prevailing tendencies in the literature rather than definitive results for every case. A consistent pattern is that KANs—particularly when equipped with specialized basis functions—tend to match or outperform vanilla MLPs and PINNs in terms of accuracy and convergence speed.
Related Surveys. A number of recent surveys have appeared on KANs, each highlighting different facets of this emerging architecture. Notable contributions include those by Andrade et al. [29], Basina et al. [30], Beatrize et al. [31], Faroughi et al. [32], Kilani et al. [33], Hou et al. [34], Dutta et al. [35], Essahraui et al. [36], and Somvanshi et al. [37], which together provide valuable entry points into the literature. Our review builds on these efforts by seeking to offer a more systematic and comprehensive perspective: rather than cataloguing studies in isolation, we integrate theoretical, architectural, optimization, and application viewpoints into a structured roadmap. By combining comparative analysis, methodological insights, and practical guidance, our goal is to provide a resource that complements existing surveys and serves readers aiming to both understand and apply KANs.
Contributions. To the best of our knowledge, this review is:
  • the first to systematically connect KAN architectures with the Kolmogorov superposition theorem;
  • the first to comprehensively bridge KANs with classical kernel methods;
  • the first to compile and compare reported accuracy, convergence behavior, and computational cost between KANs and MLPs across tasks;
  • the first to provide a comprehensive and unified overview of basis functions used in KANs;
  • the first to collect and organize a wide range of open-source KAN implementations.
  • the most complete to date provide a structured and integrative perspective on KANs by unifying theoretical foundations, architectural designs, optimization strategies, and applications into a coherent roadmap for both understanding and practice.
Organization and Key Insights. This review is structured to build a comprehensive understanding of Kolmogorov-Arnold Networks, moving from foundational theory to practical application. Section 2 establishes that KANs are not exact realizations of the Kolmogorov Superposition Theorem, but rather data-driven approximations inspired by its compositional structure, using learned smooth functions and modern deep architectures. Section 3 demystifies their behavior by showing that shallow, one-dimensional KANs are mathematically equivalent to classical kernel methods, differing only in how their coefficients are optimized. Section 4 demonstrates the formal equivalence between KANs and MLPs, highlighting that KANs achieve superior parameter efficiency and different optimization dynamics by relocating nonlinearities from nodes to edges.
Building on this, Section 5 presents the choice of basis function as a central design axis, providing a guide for selecting bases (e.g., splines, polynomials, Fourier) to match the properties of the target function (smooth, periodic, or discontinuous). Sections 6 and 7 then catalog practical techniques—such as physics-informed constraints, adaptive grids, decomposition strategies, and parallel implementations—that systematically enhance KAN accuracy and efficiency. Section 8 covers methods like sparsity, pruning, and Bayesian formulations that improve interpretability and generalization.
Finally, Section 9 explains the theoretical advantages of KANs, showing they achieve faster convergence rates, suffer less from spectral bias, and benefit from better-conditioned NTK dynamics. These insights are distilled into a practical ”Choose-Your-KAN” guide in Section 10. We conclude by identifying open research challenges in Section 11 and summarizing our findings in Section 12.
@@
Table 1: Representative performance trends comparing KAN-based models with MLPs. Results are aggregated from the literature and reflect general outcomes. “Slower training” refers to higher per-iteration cost, though total wall-clock time may differ due to fewer epochs to convergence. All rows compare KANs to plain MLP/PINN baselines without advanced features.
Ref. Accuracy Convergence / Time (per iter.) Basis Functions
@@
2 Bridging KANs and KST
Hilbert’s Question and Superpositions. At the 1900 International Congress of Mathematicians, Hilbert posed what later became known as Hilbert’s 13th Problem [78]. He asked whether every continuous function of three variables,
$f:[0,1]^{3}\to\mathbb{R},$
can be represented as a finite superposition of continuous functions of at most two variables. Hilbert conjectured that such a representation was impossible, suggesting that genuinely multivariate functions could not be decomposed into simpler lower-arity components.
To clarify what is meant by a superposition in this question—without addressing the existence of such representations—consider a formal expression of the form
@@ 2 Bridging KANs and KST
which illustrates how a function of three variables may be written as nested compositions of bivariate and univariate functions. Even the elementary identity
@@
is a simple instance of a superposition, obtained by repeated binary operations.
Contrary to Hilbert’s conjecture, the seminal results of Kolmogorov and Arnol’d showed that every continuous function of $s$ variables can indeed be represented using only sums and superpositions of univariate continuous functions. This discovery, now known as the Kolmogorov–Arnold superposition theorem, radically changed the understanding of functional representation and forms the theoretical foundation underlying modern Kolmogorov–Arnold Networks.
Kolmogorov and Arnol’d: resolving Hilbert’s 13th Problem. In a short 1956 note [79], Kolmogorov proposed the radical idea that multivariate continuous functions might be representable through superpositions of functions of fewer variables. This was not yet a theorem, but a conceptual breakthrough supported by constructive examples. In 1957, Arnol’d confirmed the conjecture for the three-variable case [80], providing an explicit representation of a general function $f(x,y,z)$ as a superposition of continuous bivariate functions. Later that same year, Kolmogorov established the first complete theorem [81], proving that any continuous function of $n>2$ variables can be expressed using only univariate functions and addition. This result definitively resolved Hilbert’s 13th Problem.
Theorem 1 (Kolmogorov Superposition Theorem (KST) [81])
Let $E_{1}=[0,1]$ and $E_{n}=[0,1]^{n}$. For every integer $n>2$, there exist continuous universal inner functions
@@
independent of the target function $f$, such that every continuous function $f\in C(E_{n})$ admits the representation
@@
where the outer functions $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ are continuous and depend on $f$, while the inner functions $\phi_{q,p}$ are universal.
The KST shows that multivariate continuity can be synthesized entirely from univariate nonlinearities and addition, a structural insight that later motivated two–layer interpretations and neural–network analogies.
@@
Refinements to KST. Kolmogorov’s constructive formulation (1) established the universality of the inner functions $\phi_{q,p}$, but it did so at considerable structural cost: the representation requires $2n^{2}{+}n$ distinct inner functions and 2n+12n{+}1 outer functions. While theoretically elegant, this level of redundancy makes the original form far too complex for computation or numerical approximation. The decades following Kolmogorov’s work therefore focused on two goals: (i) reducing the number of unique functions needed, and (ii) improving their regularity. These efforts produced the modern, compressed understanding of KST that underlies computational constructions and contemporary reinterpretations, including recent analyses by Ismailov and Ismailov [87].
(i) Reducing the number of functions. Sternfeld [89] showed that the number of summands $2n+1$ in the Kolmogorov superposition formula is optimal in a functional–analytic sense and cannot be reduced within the class of continuous superpositions. However, his functional–analytic framework shows that these 2n+12n{+}1 summands do not require 2n+12n{+}1 different inner or outer functions: the same small collection of functions can be reused repeatedly. Thus the number of distinct functions can be much smaller, even though the number of summands is fixed.
Lorentz’ reduction of outer functions. Lorentz’s 1962 exposition [84] clarified Kolmogorov’s argument and showed that the 2n+12n{+}1 outer functions $\Phi_{q}$ can be replaced by a single continuous function $\Phi$.
Theorem 2 (Kolmogorov superposition in two variables [84])
There exist continuous, monotone increasing functions
$\phi_{1,q},\,\phi_{2,q}:[0,1]\to[0,1],\qquad q=1,\ldots,5,$
such that every continuous function $f:[0,1]^{2}\to\mathbb{R}$ admits the representation
$f(x_{1},x_{2})=\sum_{q=1}^{5}\Phi\!\bigl(\phi_{1,q}(x_{1})+\phi_{2,q}(x_{2})\bigr),$ (2)
where the outer function $\Phi:[0,2]\to\mathbb{R}$ is continuous and depends on $f$.
In this formulation, the ten inner functions $\phi_{1,q}$ and $\phi_{2,q}$ arise because the two-variable case requires five summands, each involving one univariate function of $x_{1}$ and one of $x_{2}$. These inner functions are universal and independent of $f$, while the outer function $\Phi$ depends on the specific target.
More generally, for $s$ variables Lorentz obtained
@@
Sprecher’s reduction of inner functions. Sprecher [85] achieved the corresponding simplification for the inner layer. He showed that the entire family $\{\phi_{q,p}\}$ can be generated from a single monotone universal function $\psi:E_{1}\to E_{1}$, together with fixed positive weights $\{\lambda_{p}\}_{p=1}^{n}$ and uniform shifts $x\mapsto x+tq$, where $t>0$ is an arbitrarily small rational parameter.
Theorem 3 (Sprecher 1965 [85])
For each integer $N\geq 2$ there exists a monotone increasing function $\psi:E_{1}\to E_{1}$ satisfying
$\psi\in\mathrm{Lip}\!\left(\frac{\ln 2}{\ln(2N-2)}\right),\qquad\psi(0)=0,\;\psi(1)=1,$
such that for every $n\leq N$, every $f\in C(E_{n})$, and every $\delta>0$, there exists a rational $t$ with $0<t\leq\delta$ for which
$f(x_{1},\ldots,x_{n})=\sum_{q=1}^{2n+1}g\!\left(\sum_{p=1}^{n}\lambda_{p}\,\psi(x_{p}+tq)+q\right),$ (4)
where $g:\mathbb{R}\to\mathbb{R}$ is continuous, and the coefficients $\{\lambda_{p}\}_{p=1}^{n}$ together with the function $\psi$ are universal and independent of $f$.
Sprecher’s refinement thus collapses Kolmogorov’s family of inner functions to a single universal function $\psi$, and merges the family of outer functions $\{g_{q}\}$ into a single continuous $g$. In the representation (4), all significant structural complexity lies in the universal components $\psi$ and $\{\lambda_{p}\}$, while the dependence on the target function $f$ is contained solely in the choice of $g$.
@@
(ii) Improving regularity. The inner functions in Kolmogorov’s original proof are highly irregular—Hölder continuous but nowhere differentiable. A function $u:E_{1}\to\mathbb{R}$ is Hölder continuous with exponent $\alpha\in(0,1]$ if
@@
This lack of smoothness is not accidental but a fundamental requirement. Vitushkin and Henkin [90, 91, 92] proved that no universal inner family can possess $C^{1}$ (or higher) smoothness. Although $C^{1}$ regularity implies Lipschitz continuity on compact sets, differentiability itself destroys the Kolmogorov separation property required for exact superposition. As a result, smooth inner functions are incompatible with exact KST representations.
Within the theoretical development of KST, Fridman [96] showed that the universal inner functions may be chosen Lipschitz continuous, corresponding to the maximal regularity compatible with Kolmogorov-type superpositions. Sprecher’s compressed formulation [85] can be realized within this Lipschitz framework. More recently, Actor and Knepley [98, 99] formalized this obstruction by proving that $C^{1}$ inner maps cannot satisfy the Kolmogorov separation property. These results confirm that Lipschitz continuity is optimal, and that any attempt to smooth the universal inner functions necessarily breaks exact KST.
Computational KST. Algorithmic realizations of the Kolmogorov–Arnold superposition theorem may be broadly categorized as follows:
@@
  • Exact constructive KST, where both the universal inner function and the outer functions are generated by an explicit recursive procedure;
@@
  • Approximate KST (outer-only), in which the inner function is fixed and only the outer functions are approximated or learned;
@@
  • Approximate KST (inner–and–outer), where both components are approximated within a superposition-based architecture.
Exact constructions. Early digit-based constructions by Sprecher [100, 101] provided one of the first fully computable realizations of Kolmogorov’s superposition theorem. Köppen [102] later identified flaws in the original recursion, which were corrected in the rigorous formulation of Braun and Griebel [103]. Their construction guarantees continuity, monotonicity, and convergence, and shows that the universal inner function $\psi$ is not merely an abstract object, but can be generated explicitly by a finite, algorithmic recursion.
From an intuitive standpoint, the inner function $\psi$ can be viewed as a universal nonlinear reparameterization of the input. In the Kolmogorov superposition theorem, the original input lies in $0,1]^{n}$, so the parameter $n$ is the dimension of the input space. The role of $\psi$ is to encode information from this $n$-dimensional input into a single scalar variable in a universal way.
The construction of $\psi$ involves three parameters with clearly distinct roles. The dimension parameter $n$ determines how many input variables are being folded into one and therefore directly controls how complex the encoding must be. The base parameter $\gamma$ sets the scale separation of the digit-based representation. The refinement level $k$ specifies how many recursive steps are applied. Increasing $k$ refines the same basic geometric pattern, while changing $n$ modifies the structure of that pattern itself.
Figure 1 illustrates this distinction by plotting exact finite-level inner functions $\psi_{k}(x)$ for fixed $k$ and $\gamma$, comparing the cases $n=2$ and $n=4$. When $n=2$, the inner function displays many small visible steps, giving a more oscillatory appearance and stronger apparent nondifferentiability at the plotted scale. When $n=4$, the function appears flatter over longer intervals but exhibits fewer and larger transitions. This occurs because increasing $n$ rapidly suppresses the amplitude of fine-scale refinement steps, causing local irregularities to concentrate into larger but rarer features. As a result, varying $n$ provides clearer insight into how multivariate input dimension influences the geometry of the universal inner function than varying the refinement level alone.
Figure 1: Exact finite-level inner functions $\psi_{k}(x)$ generated by the constructive Kolmogorov–Köppen inner-function scheme for fixed refinement level $k$ and base parameter $\gamma=10$, comparing the cases $n=2$ and $n=4$. The case $n=2$ exhibits many small visible steps, leading to a more oscillatory appearance, whereas $n=4$ shows flatter regions interspersed with fewer but larger transitions due to stronger suppression of fine-scale increments. Left: global 2D view on $[0,1]$. Right: zoomed 2D view on $x\in[0.4,0.5]$, highlighting the contrast in local regularity.
Related constructive and guaranteed approaches include the work of Nakamura et al. [105], who developed algorithmic constructions of both the universal inner map $\psi$ and the outer correction map $\chi$, together with certified approximation error bounds. Together, these contributions clarify not only the theoretical existence of Kolmogorov-type representations, but also their explicit realizability and geometric structure.